for the function f(x)-=x^2-2x+10+8
find all values x for which f(x)=13

Answer by stanbon(75887)

for the function f(x)-=x^2-2x+10+8
find all values x for which f(x)=13
------
Solve x^2-2x+10+8 = 13
-----
x^2-2x+5 = 0
-----
x = [2+-sqrt(4-4*5)]/2
------
x = [2+-sqrt(-16)]/2
----
x = [2+-4i]/2
x = 1+2i or x = 1-2i
-----
Cheers,
Stan H.
----------